OpenStudy (anonymous):

Determine whether the series converges absolutely, converges conditionally, or diverges

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} -1^k {(k!)^3 \over (3k)!}\]

OpenStudy (anonymous):

I got (1/27) which is less 1 so converges absolutely? Can anyone confirm this answer?

OpenStudy (anonymous):

maybe converges conditionally, not sure though... calculating it..

OpenStudy (anonymous):

are you asking this\[\sum_{n=1}^{\infty} (-1)^k {(k!)^3 \over (3k)!}\]

OpenStudy (anonymous):

is it alternate series ?

OpenStudy (anonymous):

\[ \frac{(k+1)!^3}{(3k + 3)!} \frac{3k!}{k!^3} = \frac{k+1}{3!}\]If I did not make any careless mistake. This limit does not seem to converge as k -> infinity, so it abs. diverges?

OpenStudy (anonymous):

@cinar yeah it is @bmp why is the denominator 3! ? not (3k+3)(3k+2)(3k+1)?

OpenStudy (anonymous):

I missed the parenthesis there, it should be 3(k+1)!, which is 3(k+1). But I was careless with my work up there. The series appear to diverge absolutely, but do the limit more carefully than what I did.

OpenStudy (anonymous):

I guess you can use ratio test..